### v1

# ./references.bib

Abstract
A variety of physical phenomena, such as amplification, absorption, and radiation, can be effectively described using non-Hermitian operators. However, the introduction of non-uniform non-Hermiticity can lead to the formation of exceptional points in a system’s spectrum, where two or more eigenvalues become degenerate and their associated eigenvectors coallesce causing the underlying operator or matrix to become defective. Here, we explore extensions of the Clifford and quadratic $\epsilon$-pseudospectrum, previously defined for Hermitian operators, to accommodate non-Hermitian operators and matrices, including the possibility that the underlying operators may possess exceptional points in their spectra. In particular, we provide a framework for finding approximate joint eigenstates of a $d$-tuple of Hermitian operators $\bm{A}$ and non-Hermitian operators $\bm{B}$, and show that their Clifford and quadratic $\epsilon$-pseudospectra are still well-defined despite any non-normality. We prove that the non-Hermitian quadratic gap is local with respect to the probe location when there are perturbations to one or more of the underlying operators. Altogether, this framework enables the exploration of non-Hermitian physical systems’ $\epsilon$-pseudospectra, including but not limited to photonic systems where gain, loss, and radiation are prominent physical phenomena.
1 Introduction
Localized states in physical systems have a variety of important technological applications. In multidimensional systems, states that are localized to a single dimension are useful for enabling directed transport of energy or information. Moreover, zero-dimensional cavity states have uses in storing energy and information. In systems described by Hermitian Hamiltonians, finding such localized states is generally accomplished by identifying (or designing) one or more states to be isolated in the system’s spectrum, whose corresponding local density of states (LDOS) is spatially localized. Indeed, finding classes of systems that exhibit such boundary-localized states is one of the primary motivations for studying material topology [bansil_lin_das_RevModPhys.88.021004],[hasan_kane_2010_RevModPhys.82.3045],[ozawa_et_all_RevModPhys.91.015006],[xiao_chang_niu_RevModPhys.82.1959], where bulk-boundary correspondence can guarantee such localized states.
Open systems have the property that the corresponding Hamiltonian is non-Hermitian, stemming from either a loss of energy or decoherence due to the environment. As such, for an open system, the Hamiltonian $H$ can have complex valued eigenvalues whose imaginary parts correspond to the loss (or gain) of energy or information in each associate eigenstate. A canonical class of open systems are photonic systems, which can exhibit amplification, absorption, and radiation. Here, without including microscopic descriptions of the sources of amplification or absorption, or considering the infinite space into which photons can radiate, these phenomena are instead generally considered by introducing non-Hermitian terms into the photonic system’s Hamiltonian. Nevertheless, finding low-loss localized states in photonics, and open systems more broadly, is a perennial goal, as these states exhibit Purcell enhancement of any latent nonlinear effects [purcell_spontaneous_1946].
Previously, the search for system states that are approximately localized in energy and position, i.e., that are approximate joint eigenstates of both a system’s Hamiltonian and position operators, has been studied using the system’s quadratic pseudospectrum [cerjan_loring_vides_2023]. In such an approach, the idea is to choose a $(d+1)$-tuple $\bm{\lambda}=(x_{1},...,x_{d},E)$ (or by abuse of notation also written as $(\bm{x},E)$) consisting of position and energy and extracting information from the quadratic composite operator
$Q_{(\bm{x},E)}(\bm{X},H)=\sum_{i=1}^{d}(X_{i}-x_{i})^{2}+(H-E)^{2}.$ (1)
The smallest eigenvalue will then provide a unit state that is an approximate joint eigenstate of each of the matrices $X_{i}$ and $H$. In particular the ‘probe site’ $\bm{\lambda}$ in $\mathbb{R}^{n+1}$, provides the corresponding approximate eigenvalues of the matrices. This mathematical method gives rise to the quadratic $\epsilon$-pseudospectrum and is the main tool in [cerjan_loring_vides_2023]. Similar approaches based on the system’s Clifford $\epsilon$-pseudospectrum have also been used to both identify localized states in realistic photonic systems and develop a class of invariants for material topology [cerjan_loring_pho_2022, kahlil_cerjan_loring, cerjan_prl_2024]. By itself, the inclusion of non-Hermiticity does not preclude the possibility of finding localized states generally, nor boundary localized states protected by material topology see [Esaki_Sato_Hasebe_Kohmoto_PR_2011], [kawabata_shiozaki_ueda_sato_PhysRevX.9.041015_2019], [Kawabata_etal_nature_2019], and [Shen_Zhen_Fu_PRL_2018]. However, when working with a non-Hermitian systems, we can fail to find an approximately localized state when directly applying the method outlined above with the $\epsilon$-pseudospectrum. The problem is that the quadratic composite operator in Equation 1 can fail to be normal for non-Hermitian Hamiltonians $H$ and as a result could have degenerate eigenspaces. Thus the degeneracy in the eigenspaces and eigenvalues of Equation 1 could imply the quadratic gap is not small enough. Thus we could fail to find approximately localized states since the error bounds guaranteeing a good approximately localized eigenstate depends on the magnitude of the quadratic gap function.
Pseudospectral methods are being used for non-Hermitian systems, see [kiorpelidis2024scaling]. Here, we propose pseudospectral methods for predicting approximately localized states and resonances in non-Hermitian systems in both position and energy. We modify the quadratic composite operator in [cerjan_loring_vides_2023] to a generalized version of the following,
$\text{Q}_{(\bm{x},E)}\left(\bm{X},H\right)=\sum_{i}^{d}(X_{i}-x_{i})^{2}+(H-E)^{\dagger}(H-E)$ (2)
while allowing $E$ to be complex valued since $H$ being non-Hermitian gives us the possibility of complex eigenvalues. This updates definitions for a probe site as well as the $\epsilon$-pseudospectrum. We take this idea and expand it to allow for multiple non-Hermitian operators in the construction, in particular non-normal operators ($AA^{\dagger}\not=A^{\dagger}A$). We prove Lipschitz continuity of the corresponding gap functions. We also prove that under perturbations to a physical system’s Hamiltonian, the quadratic gap function allows us to find approximately localized states with small loss based on the probe site. Along the way, we investigate the Clifford composite operator and argue why the radial gap can be especially helpful in non-Hermitian systems with exceptional points. We also make a connection between the different gap functions and highlight the theory with numerical simulations. The theory and numerical simulations allow us to consider a system with an exceptional point and how the radial gap function helps spot the exceptional point. We also use the theory for numerical simulations on a line gapped system. Overall, one of our goals is to use the non-Hermitian pseudospectral methods we develop to design systems where we have some level of control over these long-lived lossy states. In particular, we anticipate that a modified quadratic $\epsilon$-pseudospectrum will have applications in designing photonic structures for enhancing light-matter interactions.
The remainder of the paper is outlined as follows: In Section 2, we introduce the idea of joint eigenstates between two matrices or observables and review known results for when all matrices involved are Hermitian. We finalize by highlighting some ideas from [cerjan_loring_vides_2023]. In Section 3, we consider a specific construction of the non-Hermitian spectral localizer in [cerjan_koekebier_Schulz_Baldes], which we call the Clifford composite operator $L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)$, and note some salient properties. We also briefly review the non-Hermitian Clifford linear gap and introduce a new gap called the non-Hermitian Clifford radial gap; the idea being that the Clifford linear gap will be more useful for line gapped physical systems, while the clifford radial gap could be more useful for point gapped systems. Additionally, we prove some results about the gaps and finalize with a conjecture. In Section 4, we motivate the construct the quadratic composite operator $\text{Q}_{(\bm{\lambda},\nu)}\left(\bm{A},\bm{B}\right)$ in Equation 2 by computing
$L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right).$
We define a quadratic gap $\mu_{(\bm{\lambda},\nu)}^{\textup{Q}}\left(\bm{A},B\right)$ and prove it’s equivalence to other numerical values. We end the section by proving Lipschitz continuity of the quadratic gap as well as provide important corollaries about joint approximate eigenstates. In Section 5, we provide error estimates between the non-Hermitian Clifford linear gap, non-Hermitian Clifford radial gap, and quadratic gap. In Section 6, we show that the quadratic gap is local, meaning that far away perturbations of $B$ relative to probe site $(\bm{\lambda},\bm{\nu})$ will have little effect on the corresponding non-Hermitian quadratic gap at that point. In Section 7, we apply the non-Hermitian Clifford linear gap, radial gap, and non-Hermitian quadratic gap functions to a two level non-Hermitian System. We argue why the non-Hermitian radial gap function should be used instead of the gap mentioned in [cerjan_koekebier_Schulz_Baldes] in the case of point gapped systems. In Section 8, we take a similar tight binding model as the one used in [cerjan_koekebier_Schulz_Baldes], and numerically compute the Clifford linear and quadratic gap functions to visualize were we might find approximately localized states in position and energy. Finally, we compare the Clifford linear gap to the quadratic gap function which highlights how closely the functions agree and where.
2 Pseudospectrum for joint approximate eigenstates
The main idea of pseudospectra and it’s techniques is to give equal treatment to each operator in a physical system. Instead of computing eigenstates for a Hamiltonian and then computing expectation values, pseudospectral techniques combine all of the incompatible observables/matrices in the system into a single composite operator. The spectrum of said composite operator give us information about states approximately localized at some position and energy when we consider a system’s position and Hamiltonian operators, see [krejcirik_siegl_tater_viola_2015_10.1063/1.4934378],[loring_schulz-baldes_2017_12504413820170101],[loring_schulz-baldes_2020_edsgcl.62678504020200301],[loring],[trefethen_1997_edsjsr.213303719970901],[trefethen_embree_2005_MR2155029],[trefethen_reddy_driscoll_1993_6dce06aa-3a1d-396a-ad07-b01838717a11]. When the matrices commute it can be show that there is a unit state that is an eigenstate (with potentially different eigenvalue) for each observable. This is a standard result in the theory of linear algebra. However it is also known that incompatible observables do not commmute, this tells us that there does not exist a joint eigenstate of the corresponding observables. The main idea however is that we are going to develop a framework in which we can find a unit eigenstate that is an approximate joint eigenstate of multiple matrices/observables through pseudospectral methods. In order to motivate the study of approximate joint eigenstates and as a result the study of $\epsilon$-pseudospectra; we begin with the study of joint eigenstates from commuting matrices/observables.
In the study of linear algebra it is known that a $d$ tuple of diagonalizable matrices $\bm{A}$ pairwise commute if and only if there exist a joint eigenstate. What this means is that if we have two matrices $A,B$, then the commutator is zero ($[A,B]=0$) then there is a state $\bm{\psi}$ such that $A\phi=\lambda_{A}\bm{\psi}$ and $B\bm{\psi}=\lambda_{B}\bm{\psi}$. To be more precise Theorem 2.3.3 in [horn_johnson] tells us that there is a unitary matrix $U$ that simultaneously transforms $A_{i}$ into upper triangular matrices, given a $d$ tuple of commuting normal matrices $\bm{A}$. In the situation were each matrix in $\bm{A}$ is normal ($A_{i}A_{i}^{\dagger}=A_{i}^{\dagger}A_{i}$) the theorem tell us that $A_{i}$ can be simultaneously diagonalized by the same unitary transformation.
If $H$ is some Hamiltonian and $X$ is a position observable, we know that in general $[X,H]\not=0$. In the special case of an atomic limit, we do have $[X,H]=0$. However, most materials are not in the atomic limit. The situation of commutators of observables not commuting ($[X,H]\not=0$) is due to a fundamental idea that can be generalized to multiple dimensions. In particular $\left[\frac{d}{dx},x\right]\not=0$, in other words as operators the derivative and position don’t commute. The physics comes into play when we note that a Hamiltonian is usually a differential operator.
Thus if we have a physical system the best we can hope for are states that are approximately localized in position and energy. In other words if we have some number of position observables $\bm{X}$, and corresponding Hermitian Hamiltonian $H$, Proposition II.1 in [cerjan_loring_vides_2023] answers the question of looking for $(\bm{x},E)$, and $\bm{\psi}$ so that
$X_{i}\bm{\psi}\approx x_{i}\bm{\psi}\mbox{ , }\hskip 7.22743ptH\bm{\psi}\approx E\bm{\psi}.$ (3)
Indeed the approach that is taken is to utilize pseudospectral methods. Broadly speaking pseudospectra is a mathematical tool to determine approximate spectrum for non-commuting matrices. To find approximate joint eigenstates of matrices or observables, the idea is to consider various eigenvalue problems $(A_{i}-\lambda_{i})\bm{\psi}$, with $A_{i}$ Hermitian. Then we take a non trivial Clifford representation $\Gamma_{i}^{\dagger}=\Gamma_{i}$, $\Gamma_{i}^{2}=I$, $\Gamma_{i}\Gamma_{k}=-\Gamma_{k}\Gamma_{i}$ for $i\not=k$ to construct the so called spectral localizer mentioned in Section I of [cerjan_loring_vides_2023] as well as Section 1 of [loring]. Given $\bm{A}$ in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ with each $A_{i}$ being Hermitian operators, and probe site $\bm{\lambda}\in\mathbb{R}^{d+1}$,
$L_{\bm{\lambda}}(\bm{A})=\sum_{i=1}^{d+1}(A_{i}-\lambda_{i})\otimes\Gamma_{i}$
is the spectral localizer.
By properties of the tensor product and the $A_{i}$ matrices, the spectral localizer is itself a Hermitian matrix. From here a so called spectral localizer gap is defined and has alternative characterizations. It is defined explicitly in Section I of [cerjan_loring_vides_2023]. However it first appears implicitly in Definition 1.1 of a so called Clifford $\epsilon$-pseudospectrum in [loring].
Note that for a $d=3$ dimensional system we could use the following as a Clifford representation
$\Gamma_{1}=\sigma_{x}\otimes\sigma_{x},\Gamma_{2}=\sigma_{y}\otimes\sigma_{x},\Gamma_{3}=\sigma_{z}\otimes\sigma_{x},\Gamma_{4}=I_{2}\otimes\sigma_{z}.$
This tells us that
$L_{(\kappa x,\kappa y,\kappa z,E)}^{\textup{3d}}\left(X,Y,Z,H\right)=\kappa(X-x)\otimes\Gamma_{1}+\kappa(Y-y)\otimes\Gamma_{2}+\kappa(Z-z)\otimes\Gamma_{3}+(H-E)\otimes\Gamma_{4}.$
We could have also used other matrices that satisfy the Clifford relations, such as
$\Gamma_{1}=\sigma_{x}\otimes\sigma_{x},\Gamma_{2}=\sigma_{y}\otimes\sigma_{x},\Gamma_{3}=\sigma_{z}\otimes\sigma_{x},\Gamma_{4}=I_{2}\otimes\sigma_{y}$
for example. In any case it is good to note that there are other possible representations. This is proven in [cerjan_loring_2024_127892] by Cerjan and Loring.
We will call $\sigma_{\min}(A)$ the smallest singular value of a matrix $A$. We then say that $\textup{spec}(A)$ is the set of eigenvalues of an operator $A$. In particular we will take the convention that the spectrum of an arbitrary matrix $A$ will be enumerated so that the eigenvalues are ordered by their real parts. We will also assume for the remainder of the paper unless otherwise stated that a matrix norm is assumed to be the spectral/operator norm. In other words if $A$ is a matrix then the assumed norm for matrices is
$\left\lVert A\right\rVert=\max_{\left\lVert\bm{\psi}\right\lVert=1}\lVert A\bm{\psi}\rVert.$
The smallest singular value of the spectral localizer plays a major role in an attempt at finding states that are localized in it’s various obeservables/operators. For this reason, in this context it is given a name. Specifically given $\bm{A}$ in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ with each $A_{i}$ being Hermitian operators, and probe site $\bm{\lambda}$ in $\mathbb{R}^{d+1}$, the spectral localizer gap is defined as
$\mu^{\text{C}}_{\bm{\lambda}}(\bm{A})=\sigma_{\min}\left(L_{\bm{\lambda}}\left(\bm{A}\right)\right).$
This gap function is historically used implicity in the definition of the Clifford $\epsilon$-pseudospectrum that is used in Section 1 of [loring]. Then as it’s signifiance grew it warranted a definition, and was defined explicitly in Section I of [cerjan_loring_vides_2023]. In particular, given $\bm{A}$ in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ with each $A_{i}$ being Hermitian operators the Clifford $\epsilon$-pseudospectrum is defined as
$\Lambda_{\epsilon}^{\text{C}}(\bm{A})=\left\{\bm{\lambda}\in\mathbb{R}^{d+1}\mid\mu_{\bm{\lambda}}^{\text{C}}\left(\bm{A}\right)\leq\epsilon\right\}$
With all this in place, one of the main results of [loring] is finding a joint approximate states with corresponding probe site each corresponding to an element in the $\epsilon$-pseudospectrum. Lemma 1.2 and 1.3 in [loring] give formal error estimates, on the quantities $\left\lVert A_{i}\bm{\psi}-\lambda_{k}\bm{\psi}\right\rVert$.
We will prove a slightly generalized version of these two Lemmas in Section 3. The main idea being that a small gap, or the pseudospectrum provides us with a recipe to get our hands on a unit state $\bm{\psi}$ that is approximately localized in position and space.
Now suppose that the operators we are working with are in $\mathcal{M}_{n}(\mathbb{C})$, for every two more operators you have in your system the dimension of the spectral localizer doubles. For 2 and 3 observables we have the spectral localizer is in $\mathcal{M}_{2n}(\mathbb{C})$, for 4 and 5 observables the spectral localizer is in $\mathcal{M}_{4n}(\mathbb{C})$, etc. This could be an issue when it comes to numerical computations as $n$ can be very large to begin with. However, the details depend on the problem at hand.
The paper [cerjan_loring_vides_2023] tackles this by considering the fact that for Hermitian matrices $A$, $A\bm{\psi}=0$ iff $A^{2}\bm{\psi}=0$. Indeed if we square the spectral localizer we get
$\left(L_{\bm{\lambda}}(\bm{A})\right)^{2}=\left(\sum_{i=1}^{d+1}(A_{i}-\lambda_{i})^{2}\otimes I\right)+\left(\sum_{i\not=k}[A_{i},A_{k}]\otimes\Gamma_{i}\Gamma_{k}\right).$
Indeed when the commutator terms are small in spectral norm, the predominant part of the squared Clifford composite operator is the first sum. This is a block diagonal matrix
$\left(\sum_{i=1}^{d+1}(A_{i}-\lambda_{i})^{2}\right)\otimes I.$ (4)
The eigenvalues of Equation 4 are determined by the repeated block matrix in 5.
$\sum_{i=1}^{d+1}(A_{i}-\lambda_{i})^{2}$ (5)
Equation (5) is taken to be the quadratic composite operator in Section 1 of [cerjan_loring_vides_2023]. After some properties are proven about the smallest eigenvalue of the quadratic composite operator, the quadratic gap function is defined.
Given $\bm{A}$ in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ where each of the $A_{i}$ are Hermitian operators and probe site $\bm{\lambda}$ in $\mathbb{R}^{d+1}$, the quadratic gap is defined to be
$\mu_{\bm{\lambda}}^{\text{Q}}(\bm{A})=\sqrt{\sigma_{\min}\left(\text{Q}_{\bm{\lambda}}(\bm{A})\right)}.$
The square root is a consequence of proving this quantity is related to the smallest singular value of a related matrix $M_{\bm{\lambda}}(\bm{A})$ which is the topic of Proposition II.1 in [cerjan_loring_vides_2023]. Note that by construction, the quadratic composite operator has real eigenvalues and is positive semidefinite. Thus the quadratic gap is well defined. Informaly a key idea with this gap function is that
$\mu_{\bm{\lambda}}^{\text{C}}(\bm{A})=\sigma_{\min}\left(\left(L_{\bm{\lambda}}\left(\bm{A}\right)\right)^{2}\right)\approx\sigma_{\min}\left(\text{Q}_{\bm{\lambda}}\left(\bm{A}\right)\right)\leq\mu_{\bm{\lambda}}^{\text{Q}}\left(\bm{A}\right)$
when $\mu_{\bm{\lambda}}^{Q}(\bm{A})\leq 1$. This also provides a corresponding state $\bm{\psi}$ that is an approximate eigenstate of the constituent observables/matrices. For a Comparison between the spectral localizer gap and quadratic gap function, see Proposition II.4 in [cerjan_loring_vides_2023]. The main idea is that the Clifford and quadratic gaps are close if the commutators $[A_{i},A_{k}]$ are small, in other words the matrices $A_{i}$ nearly pairwise commute.
3 Clifford composite operator
As mentioned in the introduction, a property of photonic systems is that they exhibit non-Hermitian Hamiltonians. This motivates the question: What happens when $A_{d+1}=B$ is non-Hermitian in the construction of the spectral localizer and in the quadratic composite operator? Or in a physical system setting, what happens when we have a non-Hermitian Hamiltonian $H$ in the construction of the corresponding composite operators? We take inspiration from the constructions and definitions in [cerjan_koekebier_Schulz_Baldes] to motivate this Section. In the process we also introduce a new gap function and explain some of it’s properties.
One option for a construction of an equivalent spectral localizer is to consider a matrix decomposition of $B$ into the Hermitian part $B_{\textup{H}}$ and it’s skew Hermitian part $B_{\textup{S}}$ shown in Equation (6)
$B=B_{\textup{H}}+B_{\textup{S}},\quad B_{\textup{H}}=(B+B^{\dagger})/2,\quad B_{\textup{S}}=-i(B-B^{\dagger})/2.$ (6)
Both $B_{\textup{H}}$ and $B_{\textup{S}}$ are Hermitian, and thus we can use these together with position observables in the Clifford composite operator, and thus we would be able to find a unit state $\bm{\psi}$ that is approximately localized in position and energy given the theory above. In particular we could construct
$L_{(\bm{x},\textrm{Re}(E),\textrm{Im}(E))}(\bm{X},B_{\textup{H}},B_{\textup{S}}).$
However this goes against the intuition that the Clifford composite operator should have similar properties as the Hamiltonian, in this case hermiticity or lack thereof.
Also, nothing necessarily stops us from using $B$ in the construction of the spectral localizer in [loring] and [cerjan_loring_vides_2023] with complex $E$. In other words $L_{(\bm{x},E)}(\bm{X},B)$. This makes it so that the resulting operator is not only not Hermitian, It is also not normal. We will be following the work in [cerjan_koekebier_Schulz_Baldes] where a special case of the non-Hermitian spectral localizer which we will refer to as the Clifford composite operator is utilized.
In the construction of the spectral localizer in Section I of [cerjan_loring_vides_2023], an arbitrary Clifford representation can be chosen. However, in the following construction for the Clifford composite operator we use a specific Clifford representation. The construction of said Clifford representation is as follows. Note that $\sigma_{x},\sigma_{y},\sigma_{z}$ are the Pauli spin matrices. For $d=2$ we use $\sigma_{x}$ and $\sigma_{y}$ and for $d=1$ we just drop $\sigma_{y}$. For $d=3$ we use $\Gamma_{1}=\sigma_{x},\Gamma_{2}=\sigma_{y},\Gamma_{3}=\sigma_{z}$. The Construction for $d=5$ is $\Gamma_{1}=\sigma_{x}\otimes\sigma_{x}$, $\Gamma_{2}=\sigma_{y}\otimes\sigma_{x}$, $\Gamma_{3}=\sigma_{z}\otimes\sigma_{x}$, $\Gamma_{4}=I_{2}\otimes\sigma_{y}$, $\Gamma_{5}=I_{2}\otimes\sigma_{z}$. Then suppose that $\{\gamma_{i}\}_{i=1}^{d}$ represents a Clifford representation for $d$ odd, we construct a Clifford representation for $d+2$ by computing
$\Gamma_{i}$ $=\gamma_{i}\otimes\sigma_{x}\text{ for all }1\leq i\leq d$
$\Gamma_{d+1}$ $=I_{m}\otimes\sigma_{y}$ (7)
$\quad\Gamma_{d+2}$ $=I_{m}\otimes\sigma_{z}$
with
$m=2^{\left\lfloor d/2\right\rfloor}.$ (8)
If we want the Clifford representation for $d$ even, we just construct the $d+2$ Clifford representation and discard $I_{m}\otimes\sigma_{y}$. We now proceed with the definition of the Clifford composite operator as follows.
Definition 3.1.
Given $\bm{A}$ in $\Pi_{i=1}^{d}\mathcal{M}_{n}(\mathbb{C})$ where each of the $A_{i}$ are Hermitian operators, a non-Hermitian matrix $B$, and a probe site $(\bm{\lambda},\nu)$ in $\mathbb{R}^{d}\oplus\mathbb{C}$ we define the Clifford composite operator to be
$L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)=\sum_{i=1}^{d}(A_{i}-\lambda_{i})\otimes\Gamma_{i}+(B-\nu)\otimes\begin{bmatrix}I_{m}&\\ &0_{m}\end{bmatrix}+(B-\nu)^{\dagger}\otimes\begin{bmatrix}0_{m}&\\ &-I_{m}\end{bmatrix}$
Where the specific matrices $\{\Gamma_{i}\}_{i=1}^{d+1}$ are constructed as in Equations 3.
Again, this is a special case of the spectral localizer in [cerjan_koekebier_Schulz_Baldes]. From here we can manually compute and determine that the Clifford composite operator is not normal. This can be seen below when we compute the product
$\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)^{\dagger}\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)$ $=\left(\sum_{i=1}^{d}(A_{i}-\lambda_{i})^{2}\otimes I\right)+\left(\sum_{i<j}\left[A_{i},A_{j}\right]\otimes\Gamma_{i}\Gamma_{j}\right)+F$ (9)
$\quad+(B-\nu)^{\dagger}(B-\nu)\otimes\begin{bmatrix}I_{m}&\\ &0_{m}\end{bmatrix}+(B-\nu)(B-\nu)^{\dagger}\otimes\begin{bmatrix}0_{m}&\\ &I_{m}\end{bmatrix},$
with
$F=\left(\sum_{i=1}^{d}(A_{i}-\lambda_{i})(B-\nu)\otimes\Gamma_{i}\begin{bmatrix}I_{m}&\\ &0\end{bmatrix}\right)+h.c.+\left(\sum_{i=1}^{d}(A_{i}-\lambda_{i})(B-\nu)^{\dagger}\otimes\Gamma_{i}\begin{bmatrix}0&\\ &-I_{m}\end{bmatrix}\right)+h.c.$ (10)
Then again we compute $\left(L_{(\bm{x},E)}\left(\bm{X},B\right)\right)\left(L_{(\bm{x},E)}\left(\bm{X},B\right)\right)^{\dagger}$ to show that
$\left[L_{(\bm{x},E)}\left(\bm{X},B\right),\left(L_{(\bm{x},E)}\left(\bm{X},B\right)\right)^{\dagger}\right]\not=0.$
Note that $h.c.$ in Equation 10 represents the Hermitian conjugate of the matrix preceding the term.
Despite the Clifford composite operator being non-normal, we do have a property that is just as useful if not better. If $B$ is close to being Hermitian then so too is the Clifford composite operator. Formally we have that for all $\epsilon>0$, $\left\lVert B-B^{\dagger}\right\rVert<\epsilon$ implies that if we label $L=L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)$
$\left\lVert L-L^{\dagger}\right\rVert$ $=\left\lVert\left((B-\nu)-(B-\nu)^{\dagger}\right)\otimes\begin{bmatrix}I_{m}&\\ &0_{m}\end{bmatrix}+\left((B-\nu)^{\dagger}-(B-\nu)\right)\otimes\begin{bmatrix}0_{m}&\\ &\!\!\!\!\!-I_{m}\end{bmatrix}\right\rVert$
$=\left\lVert\left((B-\nu)-(B-\nu)^{\dagger}\right)\otimes I_{m}\right\rVert$
$=\left\lVert(B-\nu)-(B-\nu)^{\dagger}\right\rVert$
$<\epsilon+2\textrm{Im}(\nu).$
This also tells us that the closeness of the Clifford composite operator to Hermitian is also dependent on how big of a value we probe in energy. In a way this is surprising from the point of view that if a matrix $A$ is Hermitian, then it follows that $A$ is normal, indeed $A=A^{\dagger}$ implies $A^{\dagger}A=AA=AA^{\dagger}$. Here we have shown that we can have a matrix that is non-normal but close to being Hermitian, showing the vastness of the set of complex $n$ by $n$ matrices. To define the gap function for the Clifford composite operator, we take the definition in Section 2 of [cerjan_koekebier_Schulz_Baldes] that is used while studying line gapped systems. In particular it is called the spectral localizer gap, but we will refer to it by a different name.
Definition 3.2.
Given $\bm{A}$ in $\Pi_{i=1}^{d}\mathcal{M}_{n}(\mathbb{C})$ where each of the $A_{i}$ are Hermitian operators, a non-Hermitian matrix $B$, and a probe site $(\bm{\lambda},\nu)$ in $\mathbb{R}^{d}\oplus\mathbb{C}$ we define the Clifford linear gap to be
$\bar{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}\left(\bm{A},B\right)=\min\big{|}\textrm{Re}\big{(}\textup{Spec}\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)\big{)}\big{|}.$
From a topological point of view, especially for line gapped systems, it seems to be instrumental to not look at $\min|\textup{Spec}(A)|$ but instead to flatten the spectrum and instead look at $\min|\textrm{Re}(\textup{Spec}(A))|$. When we are working with non-Hermitian Systems, we have to take into consideration both point gapped and line gapped systems. For this reason we propose the following definition for point gapped systems. By numerical computations in Section 7 we will demonstrate that the smallest singular value of the Clifford composite operator detects point gaps.
Definition 3.3.
Given $\bm{A}$ in $\Pi_{i=1}^{d}\mathcal{M}_{n}(\mathbb{C})$ where each of the $A_{i}$ are Hermitian operators, a non-Hermitian matrix $B$, and a probe site $(\bm{\lambda},\nu)$ in $\mathbb{R}^{d}\oplus\mathbb{C}$ we define the Clifford radial gap to be
$\dot{\mu}_{\bm{\lambda}}^{\text{C}}\left(\bm{A},B\right)=\sigma_{\min}\left(L_{\bm{\lambda}}\left(\bm{A},B\right)\right).$
One immediate consequence of the definition of the clifford radial gap is Lipschitz continuity of the Clifford radial gap in a generalized taxi-cab metric. Lipschitz continuity implies that the Clifford radial gap is differentiable almost everywhere. This also means that when trying to generate a grid size for numerical simulations, it can be determined ahead of time based on the level of precision desired between data points.
Proposition 3.4.
Given $\bm{A},\bm{C}\in\Pi_{i=1}^{d}\mathcal{M}_{n}(\mathbb{C})$ where each of the $A_{i}$, $C_{i}$ are Hermitian operators, non-Hermitian matrices $B$, $D$, and probe sites $(\bm{\lambda},\nu),(\bm{\alpha},\beta)\in\mathbb{R}^{d}\oplus\mathbb{C}$ then
$\left|\dot{\mu}_{\bm{\lambda}}^{\text{C}}\left(\bm{A},B\right)-\dot{\mu}_{\bm{\alpha}}^{\text{C}}\left(\bm{C},D\right)\right|\leq\left(\sum_{i=1}^{d}\left\lVert A_{i}-C_{i}\right\rVert+\left\lVert\lambda_{i}-\alpha_{i}\right\rVert\right)+\left\lVert B-D\right\rVert+\left\lVert\nu-\beta\right\rVert.$
Proof.
First by triangle inequality we have that
$\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)-L_{(\bm{\alpha},\beta)}\left(\bm{C},D\right)\right\rVert$ $\leq\sum_{i=1}^{d}\left\lVert\left((A_{i}-C_{i})+(\lambda_{i}-\alpha_{i})I_{2m}\right)\otimes\Gamma_{i}\right\rVert$
$\hskip 21.68121pt+\left\lVert\left((B-D)+(\nu-\beta)I_{m}\right)\otimes\begin{bmatrix}I_{m}&\\ &0_{m}\end{bmatrix}\right\rVert$
$\leq\left(\sum_{i=1}^{d}\left\lVert(A_{i}-C_{i})+(\lambda_{i}-\alpha_{i})I_{2m}\right\rVert\right)+\left\lVert(B-D)+(\nu-\beta)I_{m}\right\rVert$
$\leq\left(\sum_{i=1}^{d}\left\lVert A_{i}-C_{i}\right\rVert\right)+\left\lVert B-D\right\rVert+\left(\sum_{i=1}^{d}\left\lVert\lambda_{i}-\alpha_{i}\right\rVert\right)+\left\lVert\nu-\beta\right\rVert$
Lastly we make use of the lipschitz continuity of the smallest singular value of a matrix to get
$\left|\dot{\mu}_{\bm{\lambda}}^{\text{C}}\left(\bm{A},B\right)-\dot{\mu}_{\bm{\alpha}}^{\text{C}}\left(\bm{C},D\right)\right|\leq\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)-L_{(\bm{\alpha},\beta)}\left(\bm{C},D\right)\right\rVert$
The Clifford linear and radial gap functions is what we will used to define the corresponding $\epsilon$-pseudospectrum as follows.
Definition 3.5.
Given $\bm{A}$ in $\Pi_{i=1}^{d}\mathcal{M}_{n}(\mathbb{C})$ with each $A_{i}$ being Hermitian, $B\in\mathcal{M}_{n}(\mathbb{C})$ being non-Hermitian the Clifford linear $\epsilon$-pseudospectrum is defined as
$\bar{\Lambda}_{\epsilon}^{\text{C}}(\bm{A},B)=\left\{(\bm{\lambda},\nu)\in\mathbb{R}^{d}\oplus\mathbb{C}\mid\bar{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}\left(\bm{A},B\right)\leq\epsilon\right\}$
and the Clifford radial $\epsilon$-pseudospectrum defined as
$\dot{\Lambda}_{\epsilon}^{\text{C}}(\bm{A},B)=\left\{(\bm{\lambda},\nu)\in\mathbb{R}^{d}\oplus\mathbb{C}\mid\dot{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}\left(\bm{A},B\right)\leq\epsilon\right\}.$
Note that when $B$ is Hermitian and $\nu$ is real then we have
$\Lambda_{\epsilon}^{\text{C}}(\bm{A},B),=\bar{\Lambda}_{\epsilon}^{\text{C}}(\bm{A},B)=\dot{\Lambda}_{\epsilon}^{\text{C}}(\bm{A},B).$
The idea is that either the Clifford linear or radial $\epsilon$-pseudospectrum will provide a tuple $(\bm{x},E)$ and a unit eigenstate $\bm{\psi}$ so that (3) holds. As stated in Section 2 we will prove a similar formulation of Lemma 1.2 and 1.3 of [loring]. In other words the $\epsilon$-pseudospectrum gives rise to a unit state that is approximately localized in position and energy.
Proposition 3.6.
If $(\bm{\lambda},\nu)$ in $\dot{\Lambda}_{\epsilon}^{\text{C}}\left(\bm{A},B\right)$ for some $\epsilon_{1}\geq 0$ and
$\left(\sum_{i\not=k}\left\lVert[A_{i},A_{k}]\right\rVert\right)+\lVert F\rVert\leq\epsilon_{2}$
then there is a unit state $\bm{\psi}$ such that
$\sqrt{\left(\sum_{i=1}^{d}\left\lVert A_{i}\bm{\psi}-\lambda_{i}\bm{\psi}\right\rVert^{2}\right)+\left\lVert B\bm{\psi}-\nu\bm{\psi}\right\rVert^{2}}\leq\sqrt{2m}\sqrt{\epsilon_{1}^{2}+\epsilon_{2}},$
with $F$ as in Equation (10) and $m$ as in Equation (8).
Proof.
Suppose that we have selected $\Gamma_{i}$ is constructed as above so that $g=2m$ with $m$ as in Equation 8. Without loss of generality we will consider probe site $(\bm{\lambda},\nu)=(\bm{0},0)$. Suppose then that $\epsilon=\dot{\mu}_{(\bm{0},0)}^{\text{C}}\left(\bm{A},B\right)$. Since $\epsilon$ is a singular value of $L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)$, there is a corresponding right singular vector $\bm{\Psi}$. From here we have that $\bm{\Psi}$ is an eigenvector of $L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)$ with eigenvalue $\epsilon^{2}$. By equations (9),(10) and triangle inequality we have that
$\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\otimes I_{2m}\right)+C\right)\bm{\Psi}\right\rVert$ $\leq\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\bm{\Psi}\right\rVert$
$\quad+\left\lVert\left(\left(\sum_{i<j}\left[A_{i},A_{j}\right]\otimes\Gamma_{i}\Gamma_{j}\right)+F\right)\bm{\Psi}\right\rVert$
$\leq\epsilon^{2}+\left(\sum_{i\not=k}\left\lVert[A_{i},A_{k}]\right\rVert\right)+\lVert F\rVert$
with
$C=B^{\dagger}B\otimes\begin{bmatrix}I_{m}&\\ &0_{m}\end{bmatrix}+BB^{\dagger}\otimes\begin{bmatrix}0_{m}&\\ &I_{m}\end{bmatrix}.$
Now suppose that
$\bm{\Psi}=\begin{bmatrix}\bm{\Psi}_{1}\\ \vdots\\ \bm{\Psi}_{g}\end{bmatrix}.$
So that each $\bm{\Psi}_{i}$ is in $\mathbb{C}^{n}$ with $n$ being the dimension of the operators in $\bm{A}$. Then suppose that $k$ is the index maximizing $\lVert\bm{\Psi}_{k}\rVert$ so that $\lVert\bm{\Psi}_{k}\rVert\geq 1/g$. Now we will define $\bm{\psi}=\bm{\Psi}_{k}/\lVert\bm{\Psi}_{k}\rVert$. It follows that,
$\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\right)+B^{\dagger}B\right)\bm{\Psi}_{k}\right\rVert\leq\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\otimes I_{2m}\right)+C\right)\bm{\Psi}\right\rVert$
implies
$\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\right)+B^{\dagger}B\right)\bm{\psi}\right\rVert\leq g\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\otimes I_{2m}\right)+C\right)\bm{\Psi}\right\rVert.$
Next we have that
$\sum_{i=1}^{d}\left\lVert A_{i}\bm{\psi}\right\rVert^{2}+\left\lVert B\bm{\psi}\right\rVert^{2}=\left\langle\left(\left(\sum_{i=1}^{d}A_{i}^{2}\right)+B^{\dagger}B\right)\bm{\psi},\bm{\psi}\right\rangle\leq\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\right)+B^{\dagger}B\right)\bm{\psi}\right\rVert.$
This tells completes the proof. ∎
Lemma 3.7.
Given $\bm{A}$ in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ with each $A_{i}$ being Hermitian, $B$ in $\mathcal{M}_{n}(\mathbb{C})$ non-Hermitian. If
$\sqrt{\left(\sum_{i=1}^{d}\left\lVert A_{i}\bm{\psi}-\lambda_{i}\bm{\psi}\right\rVert^{2}\right)+\left\lVert B\bm{\psi}-\nu\bm{\psi}\right\rVert^{2}+\left\lVert B^{\dagger}\bm{\psi}-\overline{\nu}\bm{\psi}\right\rVert^{2}}$ $\leq\epsilon_{1}$
$\left(\sum_{i\not=k}\left\lVert[A_{i},A_{k}]\right\rVert\right)+\lVert F\rVert$ $\leq\epsilon_{2}$
for all $i$, with $F$ as in Equation 10; then $(\bm{\lambda},\nu)$ is in $\Lambda_{\epsilon}(\bm{A},B)$ with $\epsilon=\sqrt{\epsilon_{1}+\epsilon_{2}}$ or $(\bm{\lambda},\nu)$ is in $\Lambda_{\epsilon=0}(\bm{A},B)$.
Proof.
If the Clifford composite operator is singular then we are done and $(\bm{\lambda},\nu)$ is in $\Lambda_{\epsilon=0}(\bm{A},B)$. So we will assume that the Clifford composite operator is invertible. We will use linearity to reduce to the case where $(\bm{\lambda},\nu)=\bm{0}$. If $\sqrt{\sum_{i=1}^{d}\left\lVert A_{i}\bm{\psi}\right\rVert+\left\lVert B\bm{\psi}\right\rVert+\left\lVert B^{\dagger}\bm{\psi}\right\rVert}\leq\epsilon$ for some unit $\bm{\psi}$ we will define
$\bm{\Psi}=\frac{1}{\sqrt{2m}}\begin{bmatrix}\bm{\psi}\\ \vdots\\ \bm{\psi}\end{bmatrix}$
with $m$ as in Equation 8. Then by Equation 9,
$C=B^{\dagger}B\otimes\begin{bmatrix}I_{m}&\\ &0_{m}\end{bmatrix}+BB^{\dagger}\otimes\begin{bmatrix}0_{m}&\\ &I_{m}\end{bmatrix}$
and $F$ as in Equation 10, we have
$\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\otimes I_{2m}\right)+C\right)\bm{\Psi}\right\rVert\leq\frac{1}{\sqrt{2m}}\sqrt{\left(\sum_{i=1}^{d}2m\left\lVert A_{i}^{2}\bm{\psi}\right\rVert^{2}\right)+2m\left\lVert B^{\dagger}B\bm{\psi}\right\rVert^{2}+2m\left\lVert BB^{\dagger}\bm{\psi}\right\rVert^{2}}\leq\epsilon$
implies
$\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\bm{\Psi}\right\rVert$ $\leq\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\otimes I_{2m}\right)+C\right)\bm{\Psi}\right\rVert$
$\quad+\left\lVert\left(\left(\sum_{i<j}\left[A_{i},A_{j}\right]\otimes\Gamma_{i}\Gamma_{j}\right)+F\right)\bm{\Psi}\right\rVert$
$\leq\epsilon+\left(\sum_{i\not=k}\left\lVert[A_{i},A_{k}]\right\rVert\right)+\lVert F\rVert.$
Lastly note that since the Clifford composite operator is invertible,
$\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{-1}\right\rVert^{2}=\left\lVert(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right))^{-1}\right\rVert.$
This implies that
$\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{-1}\right\rVert^{-1}$ $=\sqrt{\min_{\left\lVert\mathbf{v}\right\rVert=1}\left\lVert\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)\mathbf{v}\right\rVert}$
$\leq\sqrt{\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\bm{\Psi}\right\rVert}.$
Note that in the situation where $BB^{\dagger}=B^{\dagger}B$, we can go through the proof of Lemma 3.7 and see that we can drop the term $\left\lVert B^{\dagger}\bm{\psi}-\overline{\nu}\bm{\psi}\right\rVert$ in the assumptions. A similar result and proof for the Clifford linear $\epsilon$-pseudospectrum is not clear, we leave it as a conjecture.
Conjecture 3.8.
If $(\bm{\lambda},\nu)$ in $\bar{\Lambda}_{\epsilon}^{\text{C}}\left(\bm{A},B\right)$ then there is unit state $\bm{\psi}$ so that
$X_{i}\bm{\psi}\approx x_{i}\bm{\psi}\mbox{ , }\hskip 7.22743ptH\bm{\psi}\approx E\bm{\psi}.$ (11)
The lack of a proof does not hinder us from proceeding with the generalization of the theory in [cerjan_loring_vides_2023]. The fact that for a normal matrix $A$,
$\sigma_{\min}(A)=\min|\textup{Spec}(A)|$
makes establishing a relation between the spectral localizer gap and the quadratic gap a reasonable task with not too much work. With non-normal matrices $A$ we can have that
$\sigma_{\min}(A)<\min|\textup{Spec}(A)|.$
In any case we are working with a non-normal matrix, and with luck the matrix can be nearly Hermitian or nearly normal. However we do not work with luck and thus we need a way to characterize what we mean by nearly normal. We will make use of definition and terminology presented in [ruhe] in order to characterize deviation from normality. In particular we have,
Definition 3.9.
The deviation from normality of a square matrix $A$ is defined as
$\Delta_{2}(A)=\Delta(A)=\inf\{||U||:A=Q(D+U)Q^{\dagger}\mbox{ a Schur Decomposition}\}.$
This is in some sense equivalent to considering the quantity $||AA^{\dagger}-A^{\dagger}A||$ and formal statements comparing the two quantities can be found in Theorem 2 of [elsner_paardekooper]. Also note that when the Clifford composite operator does not deviate too much from normal then we get some decent spectral variation properties.
4 Quadratic composite operator
In [cerjan_loring_vides_2023] and [loring], it was shown that when it comes to finding unit states that are approximate joint eigenstates to the obersables/operators, both the spectral localizer’s $\epsilon$-pseudospectrum and the quadratic’s $\epsilon$-pseudospectrum provide good methods in the all Hermitian observables setting. In Section 3 we have a non-Hermitian matrix/Hamiltonian in the construction of the Clifford composite operator in order to generalized the spectral localizer $\epsilon$-pseudospectrum theory. The appeal of the spectral localizer $\epsilon$-psedudospectrum in the Hermitian case and Clifford composite operator $\epsilon$-pseudospectrum in the non-Hermitian case is that is also provides information about a system’s topology see [cerjan_loring_pho_2022],[cerjan_loring_2022],[kahlil_cerjan_loring] for the Hermitian setting and [cerjan_koekebier_Schulz_Baldes] for the non-Hermitian setting. The quadratic $\epsilon$-pseudospectrum seems to be more helpful with finding localized states as described, with some added benefits. The construction of the quadratic composite operator is inspired by Cerjan, Loring and Vides in [cerjan_loring_vides_2023], where it is defined for systems described by Hermitian Hamiltonians. The quadratic composite operator first appears implicitly in [loring]. Because of the assumption taken in these prior studies that the constituent operators in the construction are Hermitian and the ‘probe site’, i.e., the choice of $(\bm{x},E)$ is real, it is guaranteed that the corresponding $\epsilon$-pseudospectrum is real i.e. $\bm{\lambda}$ in $\mathbb{R}^{n+1}$.
One key idea is that for $A$ Hermitian, $A\bm{\Psi}=0$ if and only if $A^{2}\bm{\Psi}=0$. Thus, it is reasonable to consider a similar procedure with the Clifford composite operator in [cerjan_koekebier_Schulz_Baldes]. In particular there is reason to believe that $A^{\dagger}A\bm{\Psi}=0$ can provide useful information about $A\bm{\Psi}=0$. We know that the singular values of the $L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)$ are the square roots of the eigenvalues of
$\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)^{\dagger}\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right).$
In particular this is useful since a small deviation from normality of the Clifford composite operator implies
$\dot{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}(\bm{A},B)\approx\min\left|\textup{spec}\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)\right|$
and small deviation from being Hermitian of the Clifford composite operator implies
$\bar{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}(\bm{A},B)\approx\min\left|\textup{spec}\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)\right|$
with formal estimates proven in Section 5.
Thus we have that we can approximate the Clifford linear and radial gaps by the smallest singular value of the Clifford composite operator. By Equation 9 we have a way of defining a quadratic composite operator for the Clifford composite operator similar to the Hermitian setting in [cerjan_loring_vides_2023]. We do this by ignoring $F$ and the commutator terms in the computation of Equation 9. Specifically we could use
$\left(\sum_{i=1}^{d}(A_{i}-\lambda_{i})^{2}\otimes I\right)+(B-\nu)^{\dagger}(B-\nu)\otimes\begin{bmatrix}I_{m}&\\ &0_{m}\end{bmatrix}+(B-\nu)(B-\nu)^{\dagger}\otimes\begin{bmatrix}0_{m}&\\ &I_{m}\end{bmatrix}.$
Since the spectrum is determined by the diagonal block entries we could take one of them. Namely,
$\sum_{i=1}^{d}(A_{i}-\lambda_{i})^{2}+(B-\nu)^{\dagger}(B-\nu).$
Note we don’t consider
$\sum_{i=1}^{d}(A_{i}-\lambda_{i})^{2}+(B-\nu)(B-\nu)^{\dagger}.$
as the spectrum of both are the same with the exception of a zero eigenvalue. Instead of taking this as the definition of a quadratic composite operator, we will generalize this construction. We will allow various non-normal operators in the construction besides $B-\nu$. The proof that is to come about approximately localized eigenstates will follow with this more general construction.
Definition 4.1.
Suppose that we have tuples of matrices $\bm{A}$ in $\Pi_{i=1}^{d_{1}}M_{n}(\mathbb{C})$, $\bm{B}$ in $\Pi_{i=1}^{d_{2}}M_{n}(\mathbb{C})$ with $A_{i}$ Hermitian, $B_{j}$ non-Hermitian, and probe site $(\bm{\lambda},\bm{\nu})$ in $\mathbb{R}^{d_{1}}\oplus\mathbb{C}^{d_{2}}$ then we define the quadratic composite operator as
$\text{Q}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)=\left(\sum_{i=1}^{d_{1}}(A_{i}-\lambda_{i})^{2}\right)+\left(\sum_{j=1}^{d_{2}}(B_{j}-\nu_{j})^{\dagger}(B_{j}-\nu_{j})\right).$
Using this definition we have generalized the quadratic composite operator in [cerjan_loring_vides_2023]. In fact the following proposition generalizes Proposition II.1. in [cerjan_loring_vides_2023].
If we recall that a unit state $\bm{\psi}$ in a Hilbert space determines a probability distribution with respect to a normal matrix $M$. In particular for Hermitian matrices $M$ it is well established that we need only a complete basis for the eigenspace of $M$ to define a probablity distribution and as a result to define expectation and variance, see Appendix A of [hirvensalo_2001]. While we do not need to use the probability distribution directly, we do need the corresponding expectation and square variance for both Hermitian and non-Hermitian matrices. This allows us to talk loosely about approximate localization of a state in position, energy, momentum, frequency, etc. The mathematical notation that we will use for expectation and variance are as follows,
$\textup{E}_{\bm{\psi}}[B]$ $=\langle B\bm{\psi},\bm{\psi}\rangle$
$\textup{V}^{2}_{\bm{\psi}}[B]$ $=\langle B^{\dagger}B\bm{\psi},\bm{\psi}\rangle-\left|\langle B\bm{\psi},\bm{\psi}\rangle\right|^{2}.$
Indeed when we have that $B=A$ is Hermitian then this reduces to the usual expectation and variance of a state $\bm{\psi}$ with respect to the appropriate observable/matrix. With that in place we proceed to prove some results that link minimization of expectation and variance of matrices/observables to the quadratic composite operator.
Proposition 4.2.
Suppose that we have tuples of matrices $\bm{A}\in\Pi_{i=1}^{d_{1}}M_{n}(\mathbb{C}),\bm{B}\in\Pi_{i=1}^{d_{2}}M_{n}(\mathbb{C})$ with $A_{i}$ Hermitian, $B_{j}$ non-Hermitian, and $(\bm{\lambda},\bm{\nu})$ in $\mathbb{R}^{d_{1}}\oplus\mathbb{C}^{d_{2}}$, the following are equal:
  1. 1.
    The quantity
    $\min_{||\bm{\psi}||=1}\sqrt{\sum_{i=1}^{d_{1}}||A_{i}\bm{\psi}-\lambda_{i}\bm{\psi}||^{2}+\sum_{j=1}^{d_{2}}||B_{j}\bm{\psi}-\nu_{j}\bm{\psi}||^{2}}.$
  2. 2.
    The quantity
    $\min_{||\bm{\psi}||=1}\sqrt{\sum_{i=1}^{d_{1}}\textup{V}^{2}_{\bm{\psi}}[A_{i}]+\left(\textup{E}_{\bm{\psi}}[A_{i}]-\lambda_{i}\right)^{2}+\sum_{j=1}^{d_{2}}\textup{V}^{2}_{\bm{\psi}}[B_{j}]+\left|\textup{E}_{\bm{\psi}}[B_{j}]-\nu_{j}\right|^{2}},$
    if $B_{i}$ is normal for all $1\leq i\leq d_{2}$.
  3. 3.
    The smallest singular value of
    $M_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)=\begin{bmatrix}A_{1}-\lambda_{1}\\ \vdots\\ A_{d_{1}}-\lambda_{d_{1}}\\ B_{1}-\nu_{1}\\ \vdots\\ B_{d_{2}}-\nu_{d_{2}}\end{bmatrix}.$
  4. 4.
    The square root for the smallest eigenvalue of $\text{Q}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)$.
Proof.
To show (1) $\Leftrightarrow$ (2) we have that $A$ is Hermitian and $\bm{\lambda}\in\mathbb{R}^{d_{1}}$ we have that $(A-\lambda_{i})^{2}=(A-\lambda_{i})^{\dagger}(A-\lambda_{i})$ and thus it is enough to prove the statement for $Q_{(\vec{\bm{0}},\bm{\nu})}(\bm{0},B)$.
$||B_{j}\bm{\psi}-\nu_{j}\bm{\psi}||^{2}$ $=\langle B_{j}\bm{\psi}-\nu_{j},B_{j}\bm{\psi}-\nu_{j}\rangle$
$=\langle B_{j}\bm{\psi},B_{j}\bm{\psi}\rangle-\overline{\nu_{j}}\langle B_{j}\bm{\psi},\bm{\psi}\rangle-\nu_{j}\langle\bm{\psi},B_{j}\bm{\psi}\rangle+|\nu_{j}|^{2}\langle\bm{\psi},\bm{\psi}\rangle$
$=\left(\langle B_{j}^{\dagger}B_{j}\bm{\psi},\bm{\psi}\rangle-\left|\langle B_{j}\bm{\psi},\bm{\psi}\rangle\right|^{2}\right)+\left(\left|\langle B_{j}\bm{\psi},\bm{\psi}\rangle\right|^{2}-\overline{\nu_{j}}\langle B_{j}\bm{\psi},\bm{\psi}\rangle-\nu_{j}\langle\bm{\psi},B_{j}\bm{\psi}\rangle+|\nu_{j}|^{2}\right)$
$=\left(\langle B_{j}^{\dagger}B_{j}\bm{\psi},\bm{\psi}\rangle-\left|\langle B_{j}\bm{\psi},\bm{\psi}\rangle\right|^{2}\right)+\left|\langle B_{j}\bm{\psi},\bm{\psi}\rangle-\nu_{j}\right|^{2}$
$=\textup{V}^{2}_{\bm{\psi}}[B_{j}]+\left|\textup{E}_{\bm{\psi}}[B_{j}]-\nu_{j}\right|^{2}.$
To show (3) $\Leftrightarrow$ (4) we note that for any matrix $A$, the singular values of $A$ are the square roots of the eigenvalues of $A^{\dagger}A$. We combine this with the computation
$\left(M_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)^{\dagger}\left(M_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)=\text{Q}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right).$
To show (1) $\Leftrightarrow$ (3) we note by [golub_vanloan]
$\sigma_{\min}\left(M_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)$ $=\min_{||\bm{\psi}||=1}\left\lVert M_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\bm{\psi}\right\rVert$
$=\min_{||\bm{\psi}||=1}\left\lVert\begin{bmatrix}A_{1}\bm{\psi}-\lambda_{1}\bm{\psi}\\ \vdots\\ A_{d_{1}}\bm{\psi}-\lambda_{d_{1}}\bm{\psi}\\ B_{1}\bm{\psi}-\nu_{1}\bm{\psi}\\ \vdots\\ B_{d_{2}}\bm{\psi}-\nu_{d_{2}}\bm{\psi}\end{bmatrix}\right\rVert$
$=\min_{||\bm{\psi}||=1}\sqrt{\sum_{i=1}^{d_{1}}||A_{i}\bm{\psi}-\lambda_{i}\bm{\psi}||^{2}+\sum_{j=1}^{d_{2}}||B_{j}\bm{\psi}-\nu_{j}\bm{\psi}||^{2}}.$
Thus if we want a state that is approximately localized in position and energy we have reason to study the smallest eigenvalue of the quadratic composite operator.
Definition 4.3.
Suppose that $\bm{A}$ in $\Pi_{i=1}^{d_{1}}\mathcal{M}_{n}(\mathbb{C})$, $\bm{B}$ in $\Pi_{i=1}^{d_{2}}\mathcal{M}_{n}(\mathbb{C})$ with $A_{i}$ Hermitian, $B_{j}$ non-Hermitian and probe site $(\bm{\lambda},\bm{\nu})$ in $\mathbb{R}^{d_{1}}\oplus\mathbb{C}^{d_{2}}$, then the quadratic gap is defined as
$\mu_{(\bm{\lambda},\bm{\nu})}^{\text{Q}}\left(\bm{A},\bm{B}\right)=\sqrt{\sigma_{\min}\left(\text{Q}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)}.$
From here we define the quadratic $\epsilon$-pseudospectrum.
Definition 4.4.
Suppose that $\bm{A}$ in $\Pi_{i=1}^{d_{1}}\mathcal{M}_{n}(\mathbb{C})$, $\bm{B}$ in $\Pi_{i=1}^{d_{2}}\mathcal{M}_{n}(\mathbb{C})$ with $A_{i}$ Hermitian, $B_{j}$ non-Hermitian, and some $\epsilon\geq 0$ we define the non-Hermitian $\epsilon$-pseudospectrum as
$\Lambda_{\epsilon}^{\text{Q}}\left(\bm{A},\bm{B}\right)=\left\{(\bm{\lambda},\bm{\nu})\in\mathbb{R}^{d_{1}}\oplus\mathbb{C}^{d_{2}}\mid\mu_{(\bm{\lambda},\bm{\nu})}^{\text{Q}}\left(\bm{A},\bm{B}\right)\leq\epsilon\right\}.$
If we have $\epsilon=0$ in the definition of the $\epsilon$-pseudospectrum, there is a joint eigenstate $\bm{\psi}$ of $A_{i}$, and $B_{j}$ for all $i$ and $j$. This is more of a mathematical statement as we know that in physical systems, observables/operators do not commute and thus do not have a joint eigenspace. From this point of view, the Proposition 4.2 is a more useful result for physical systems when $\epsilon>0$.
In the study of physical systems we would like to see how the quadratic gap function behaves as we perturb our system as well as when we perturb the probe site. When we treat the quadratic gap as a function of the probe site alone, we can think of the probe site as taking a guess at a point in the approximate joint spectrum of the matrices $A_{i}$ and $B_{i}$. When we treat the quadratic gap as a function of the input matrices alone, we can think of studying perturbations to physical systems such as introducting point defects, line defects, etc. Thus, a useful property of the quadratic gap is lipschitz continuity in $((\bm{A},\bm{B}),(\bm{\lambda},\bm{\nu}))$. For one this implies that the quadratic gap function is differentiable almost everwhere in the measure theory sense. It also means that if we are running numerical simulations, determining a grid size for the probe sites can be predetermined ahead of time. In the case where only the probe site varies, machine precision $\epsilon_{\text{mach}}$ implies that between grid spaces, the precision is also $\epsilon_{\text{mach}}$.
Proposition 4.5.
Suppose that $\bm{A},\bm{C}$ in $\Pi_{i=1}^{d_{1}}\mathcal{M}_{n}(\mathbb{C}),\bm{B},\bm{D}$ in $\Pi_{i=1}^{d_{2}}\mathcal{M}_{n}(\mathbb{C})$ with $A_{i},C_{i}$ Hermitian, $B_{j},D_{j}$ non-Hermitian, and $(\bm{\lambda},\bm{\nu}),(\bm{\alpha},\bm{\beta})$ in $\mathbb{R}^{d_{1}}\oplus\mathbb{C}^{d_{2}}$, then
$\left|\mu_{(\bm{\lambda},\bm{\nu})}^{\text{Q}}\left(\bm{A},\bm{B}\right)-\mu_{(\bm{\alpha},\bm{\beta})}^{\text{Q}}\left(\bm{C},\bm{D}\right)\right|\leq\left\lVert(\bm{A},\bm{B})-(\bm{C},\bm{D})\right\rVert+\left\lVert(\bm{\lambda},\bm{\nu})-(\bm{\alpha},\bm{\beta})\right\rVert.$
Proof.
First note that
$\left\lVert\begin{bmatrix}(\lambda_{1}-\alpha_{1})I\\ \vdots\\ (\lambda_{d_{1}}-\alpha_{d_{1}})I\\ (\nu_{1}-\beta_{1})I\\ \vdots\\ (\nu_{d_{2}}-\beta_{d_{2}})I\end{bmatrix}\right\rVert^{2}$ $=\left\lVert\sum_{i=1}^{d_{1}}\left|\lambda_{i}-\alpha_{i}\right|^{2}I+\sum_{j=1}^{d_{2}}\left|\nu_{j}-\beta_{j}\right|^{2}I\right\rVert$
$=\sum_{i=1}^{d_{1}}\left(\lambda_{i}-\alpha_{i}\right)^{2}+\sum_{j=1}^{d_{2}}\left|\nu_{j}-\beta_{j}\right|^{2}$
$=||(\bm{\lambda},\bm{\nu})-(\bm{\alpha},\bm{\beta})||^{2}.$
Then by a similar argument
$\left\lVert\begin{bmatrix}(A_{1}-C_{1})\\ \vdots\\ (A_{d_{1}}-C_{d_{1}})\\ (B_{1}-D_{1})\\ \vdots\\ (B_{d_{2}}-D_{d_{2}})\end{bmatrix}\right\rVert^{2}$ $=\left\lVert\sum_{i=1}^{d_{1}}\left(A_{i}-C_{i}\right)^{2}+\sum_{j=1}^{d_{2}}\left(B_{i}-D_{i}\right)^{\dagger}\left(B_{i}-D_{i}\right)\right\rVert$
$=\sum_{i=1}^{d_{1}}\left\lVert A_{i}-C_{i}\right\rVert^{2}+\sum_{j=1}^{d_{2}}\left\lVert B_{i}-D_{i}\right\rVert^{2}$
$=||(\bm{A},\bm{B})-(\bm{C},\bm{D})||^{2}.$
It then follows that,
$\left\lVert M_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)-M_{(\bm{\alpha},\bm{\beta})}\left(\bm{C},\bm{D}\right)\right\rVert$ $\leq\left\lVert\begin{bmatrix}(A_{1}-C_{1})\\ \vdots\\ (A_{d_{1}}-C_{d_{1}})\\ (B_{1}-D_{1})\\ \vdots\\ (B_{d_{2}}-D_{d_{2}})\end{bmatrix}\right\rVert+\left\lVert\begin{bmatrix}(\lambda_{1}-\alpha_{1})I\\ \vdots\\ (\lambda_{d_{1}}-\alpha_{d_{1}})I\\ (\nu_{1}-\beta_{1})I\\ \vdots\\ (\nu_{d_{2}}-\beta_{d_{2}})I\end{bmatrix}\right\rVert$
$\leq||(\bm{A},\bm{B})-(\bm{C},\bm{D})||+||(\bm{\lambda},\bm{\nu})-(\bm{\alpha},\bm{\beta})||.$
Lastly Proposition 4.2 tells us
$\left|\mu_{(\bm{\lambda},\bm{\nu})}^{\text{Q}}\left(\bm{A},\bm{B}\right)-\mu_{(\bm{\alpha},\bm{\beta})}^{\text{Q}}\left(\bm{E},\bm{F}\right)\right|$ $=\left|\sigma_{\min}\left(M_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)-\sigma_{\min}\left(M_{(\bm{\alpha},\bm{\beta})}\left(\bm{E},\bm{F}\right)\right)\right|.$
Combine this with perturbation of singular values on page 78 of [bhatia] and we are done. ∎
5 Relationship between Clifford linear, radial and quadratic gaps
As mentioned in Section 3 we noted that the Clifford composite operator is non-normal. This means that the variation in the spectrum can vary drastically.
Despite this we have that Theorem 1 in [ruhe],
$\left|\sigma_{\min}\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)-\min\left|\textup{spec}\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)\right|\right|\leq\Delta\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right).$ (12)
By claim (ii) in [kahan], equivalence of norms and by a geometric argument we have that
$\left|\min\left|\textrm{Re}\left(\textup{Spec}\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)\right)\right|-\min\left|\textup{Spec}\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)\right|\right|\leq\sqrt{n}\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)-L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}\right\rVert$ (13)
where $n$ is the dimension of $L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)$.
Lastly
$\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)-L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}\right\rVert=\left\lVert(B-\nu)-(B-\nu)^{\dagger}\otimes I_{m}\right\rVert=\left\lVert(B-\nu)-(B-\nu)^{\dagger}\right\rVert.$ (14)
this will establish a relationship between the Clifford linear and radial gaps.
Proposition 5.1.
Suppose that we have tuples of matrices $\bm{A}$ in $\Pi_{i=1}^{d_{1}}M_{n}(\mathbb{C}),B$ in $M_{n}(\mathbb{C})$ with $A_{i}$ Hermitian, $B$ non-Hermitian, and $(\bm{\lambda},\nu)$ in $\mathbb{R}^{d_{1}}\oplus\mathbb{C}$, then
$\left|\bar{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}\left(\bm{A},B\right)-\dot{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}\left(\bm{A},B\right)\right|\leq\sqrt{n}\left\lVert(B-\nu)-(B-\nu)^{\dagger}\right\rVert+\Delta(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)),$
where $n$ is the dimension of $L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)$.
Proof.
This follows from Equations (12), (13) and (14). ∎
Indeed this formally tells us that if $\nu$ is real valued and $B$ is Hermitian, the two gaps are the same. In fact $B$ Hermitian and $\nu$ real means we have equality with the spectral localizer gap in [cerjan_loring_vides_2023], and [loring].
Proposition 5.2.
Suppose that we have tuples of matrices $\bm{A}\in\Pi_{i=1}^{d_{1}}M_{n}(\mathbb{C}),B$ in $M_{n}(\mathbb{C})$ with $A_{i}$ Hermitian, $B$ non-Hermitian, and $(\bm{\lambda},\nu)$ in $\mathbb{R}^{d_{1}}\oplus\mathbb{C}$, then
$\left|\dot{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}\left(\bm{A},B\right)-\mu_{(\bm{\lambda},\nu)}^{\text{Q}}\left(\bm{A},B\right)\right|\leq\sqrt{\sum_{i<j}||[A_{i},A_{j}]||+||F||},$
with $F$ as in equation 10.
Proof.
Note that tensoring by the identity does not affect the spectrum of a matrix (ignoring multiplicity). Therefore, given
$K=\sum_{i=1}^{d}(A_{i}-\lambda_{i})^{2}\otimes I+(B-\nu)^{\dagger}(B-\nu)\otimes\begin{bmatrix}I_{m}&\\ &0_{m}\end{bmatrix}+(B-\nu)(B-\nu)^{\dagger}\otimes\begin{bmatrix}0_{m}&\\ &I_{m}\end{bmatrix}$
we get that
$\sqrt{\sigma_{\min}\left(\text{Q}_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)}=\sqrt{\sigma_{\min}\left(K\right)}.$
We also have that
$\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)-K\right\rVert$ $=\left\lVert\sum_{i<j}[A_{i},A_{j}]\otimes\Gamma_{i}\Gamma_{j}+F\right\rVert$
$\leq\left\lVert\sum_{i<j}[A_{i},A_{j}]\otimes\Gamma_{i}\Gamma_{j}\right\rVert+\left\lVert F\right\rVert$
$\leq\sum_{i<j}\left\lVert[A_{i},A_{j}]\right\rVert+\left\lVert F\right\rVert.$
Next we have that for any positive quantities $a,b$ that $\left|\sqrt{a}-\sqrt{b}\right|\leq\sqrt{|a-b|}$. Combining this with Lipschitz continuity of singular values we get that
$\left|\dot{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}\left(\bm{A},B\right)-\mu_{(\bm{\lambda},\nu)}^{\text{Q}}\left(\bm{A},B\right)\right|$ $=\left|\sigma_{\min}(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right))-\sqrt{\sigma_{\min}(K)}\right|$
$=\left|\sqrt{\sigma_{\min}\left((L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right))^{\dagger}(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right))\right)}-\sqrt{\sigma_{\min}(K)}\right|$
$=\sqrt{\left|\sigma_{\min}\left((L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right))^{\dagger}(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right))\right)-\sigma_{\min}(K)\right|}$
$\leq\sqrt{\sum_{i<j}\left\lVert[A_{i},A_{j}]\right\rVert+\left\lVert F\right\rVert}.$
Combining the results from 5.1 and 5.2 we have the following proposition.
Proposition 5.3.
Suppose that we have tuples of matrices $\bm{A}$ in $\Pi_{i=1}^{d_{1}}M_{n}(\mathbb{C}),B$ in $M_{n}(\mathbb{C})$ with $A_{i}$ Hermitian, $B$ non-Hermitian, and $(\bm{\lambda},\nu)$ in $\mathbb{R}^{d_{1}}\oplus\mathbb{C}$, then
$\left|\bar{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}\left(\bm{A},B\right)-\mu_{(\bm{\lambda},\nu)}^{\text{Q}}\left(\bm{A},B\right)\right|$ $\leq\sqrt{n}\left\lVert(B-\nu)-(B-\nu)^{\dagger}\right\rVert+\Delta\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)+\sqrt{\sum_{i<j}\left\lVert[A_{i},A_{j}]\right\rVert+\left\lVert F\right\rVert},$
where $n$ is the dimension of $L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)$ and $F$ as in equation 10.
Proof.
By Proposition 5.1 and 5.2 we have that,
$\left|\bar{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}\left(\bm{A},B\right)-\sigma_{\min}(K)\right|$ $\leq\left|\bar{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}\left(\bm{A},B\right)-\sigma_{\min}\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)\right|+\left|\sigma_{\min}\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)-\sigma_{\min}(K)\right|$
$=\left|\bar{\mu}_{(\bm{\lambda},\bm{\nu})}^{\text{C}}\left(\bm{A},B\right)-\dot{\mu}_{(\bm{\lambda},\bm{\nu})}^{\text{C}}\left(\bm{A},B\right)\right|+\left|\dot{\mu}_{(\bm{\lambda},\bm{\nu})}^{\text{C}}\left(\bm{A},B\right)-\mu_{(\bm{\lambda},\bm{\nu})}^{\text{Q}}\left(\bm{A},B\right)\right|$
$\leq\sqrt{n}\left\lVert(B-\nu)-(B-\nu)^{\dagger}\right\rVert+\Delta\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)+\sqrt{\sum_{i<j}\left\lVert[A_{i},A_{j}]\right\rVert+\left\lVert F\right\rVert}.$
Proposition 5.2 is also equivalent Proposition II.4 in [cerjan_loring_vides_2023] when the the operator $B$ is Hermitian and the probe site $\nu$ for be is restricted to be real valued.
6 Local nature of the quadratic composite operator
The idea here with the non-Hermitian composite operator being local is that if we had a physical system, a perturbation to the Hamiltonian far away from the probe site $(\bm{\lambda},\bm{\nu})$ will have little to no effect on the quadratic gap. In other words small defects away from a probe site should not have significantly large effects when probing away from the defect. The following theorem establishes this idea. In fact this is a generalization of Theorem III.1 from [cerjan_loring_vides_2023], and the proof in in the same spirit.
For the remainder of the section we will assume that the entries of the matrix $d_{1}$ tuple in $\bm{A}$ commute and that each of the $A_{i}$ is invertible. In a physical setting where $A_{i}=X_{i}$ are position observables this is easily accomplished by shifting the system in position so that the corresponding operator is invertible. Because of the linear properties of the quadratic operator we consider $(\bm{\lambda},\bm{\nu})=\vec{\bm{0}}$. This allows us to consider the following construction. It can be though of as the Euclidian distance from the origin in spacial coordinates. Specifically,
$Z=\sqrt{A_{1}^{2}+\cdots+A_{d_{1}}^{2}}$
and
$Y=\sqrt{B_{1}^{\dagger}B_{1}+\cdots+B_{d_{2}}^{\dagger}B_{d_{2}}}.$
Theorem 6.1.
Assume that $\bm{A}$ in $\Pi_{i=1}^{d_{1}}\mathcal{M}_{n}(\mathbb{C})$ is a $d_{1}$-tuple of Hermitian matrices, $\bm{B},\bm{C}$ in $\Pi_{i=1}^{d_{2}}\mathcal{M}_{n}(\mathbb{C})$ are $d_{2}$-tuples of non-Hermitian matrices then we have that if
$\left\lVert Z^{-1}\left(\sum_{j=1}^{d_{2}}\left(B_{j}^{\dagger}C_{i}+C_{i}^{\dagger}B_{j}+C_{i}^{\dagger}C_{i}\right)\right)Z^{-1}\right\rVert\leq C$
for some $C<1$, then
$(1-C)^{1/2}\mu_{\vec{\bm{0}}}^{\text{Q}}\left(\bm{A},\bm{B}\right)\leq\mu_{\vec{\bm{0}}}^{\text{Q}}\left(\bm{A},\bm{B}+\bm{C}\right)\leq(1+C)^{1/2}\mu_{\vec{\bm{0}}}^{\text{Q}}\left(\bm{A},\bm{B}\right).$
Proof.
After direct computation we have that
$-CZ^{2}\leq\sum_{j=1}^{d_{2}}\left(B_{j}^{\dagger}C_{i}+C_{i}^{\dagger}B_{j}+C_{i}^{\dagger}C_{i}\right)\leq CZ^{2}.$
Thus,
$\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}+\bm{C}\right)$ $\leq Z^{2}+Y^{2}+CZ^{2}$
$\leq(1+C)\left(Z^{2}+Y^{2}\right)$
$=(1+C)\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}\right)$
$\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}+\bm{C}\right)$ $\geq Z^{2}+Y^{2}-CZ^{2}$
$\geq(1-C)\left(Z^{2}+Y^{2}\right)$
$=(1-C)\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}\right).$
From here we have that
$(1+C)^{-1}\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}\right)^{-1}\leq\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}+\bm{C}\right)^{-1}\leq(1-C)^{-1}\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}\right)^{-1}.$
Then since the quadratic composite operator is Hermitian Positive Semidefinite we have that
$(1+C)^{-1}\left\lVert\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}\right)^{-1}\right\rVert\leq\left\lVert\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}+\bm{C}\right)^{-1}\right\rVert\leq(1-C)^{-1}\left\lVert\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}\right)^{-1}\right\rVert.$
Finally,
$(1-C)^{\frac{1}{2}}\left\lVert\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}\right)^{-1}\right\rVert^{-\frac{1}{2}}\leq\left\lVert\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}+\bm{C}\right)^{-1}\right\rVert^{-\frac{1}{2}}\leq(1+C)^{\frac{1}{2}}\left\lVert\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}\right)^{-1}\right\rVert^{-\frac{1}{2}}.$
Note that in the proof we make extensive use that for positive semidefinite matrices $A,B$, $A\geq B$ if and only if $A-B$ is positive semidefinite. This generates an ordering on the set of positive semidefinite matrices.
7 Applications to a point gaped system
In order to encounter non-Hermiticity in physical systems one need not go very far. Indeed we consider a two level system such as the one mentioned in [tls_paper]. The two level system gives rise to a non-Hermitian Hamiltonian operator. In particular we have a point gap system. We will provide some numerical justification for why the Clifford radial gap might be the best one to use over the established Clifford linear gap. To illustrate, we first consider a general case with shifted energy as follows:
$X=\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}\mbox{ and }H=\begin{bmatrix}\Delta E+i\Delta\gamma&c\\ c&0\end{bmatrix}$
with $\Delta E,\Delta\gamma,c\in\mathbb{R}$. By doing a traditional eigenenergy and eigenstate analysis we have that the eigenenergy and eigenstate pair are
$\lambda_{\pm}(H)=\frac{(\Delta E+i\Delta\gamma)\pm\sqrt{(\Delta E+i\Delta\gamma)^{2}+4c^{2}}}{2}\hskip 7.22743pt,\hskip 7.22743pt\bm{\psi}_{\pm}=\begin{bmatrix}\lambda_{\pm}(H)/c&1\end{bmatrix}^{T}$
If we consider the case when $\Delta E=0$ then we have an exceptional point when $c=\mp\Delta\gamma/2$. In this scenario we have the eigenenergy and eigenstate pair
$\lambda_{\pm}=\frac{i\Delta\gamma}{2}\hskip 7.22743pt,\hskip 7.22743pt\bm{\psi}=\begin{bmatrix}i\Delta\gamma/c&1\end{bmatrix}^{T}$
In other words the eigenspace of $H$ becomes degenerate! The quadratic gap for this setup can get complicated, so we compute one of the exceptional point cases when $21=c=\Delta\gamma/$ 2. We will fix the position probe site of $Q$ to be $x=0$ so that the non-Hermitian Clifford and quadratic composite operators become
$L_{(0,E)}(X,H)=\begin{bmatrix}2i-E&1&-1&0\\ 1&-E&0&1\\ -1&0&-2i+\overline{E}&-1\\ 0&1&-1&-\overline{E}\\ \end{bmatrix},\mbox{ }Q_{(0,E)}(X,H)=\begin{bmatrix}\left|E\right|^{2}+4\textrm{Im}(E)+6&-2\textrm{Re}(E)-2i\\ -2\textrm{Re}(E)+2i&2+\left|E\right|^{2}\end{bmatrix}$
In Table 1 we compute the Clifford linear gap, radial gap, the quadratic gap, as well as the differences between each gap. We can see that the Clifford radial gap (Figure 2) detects the exceptional point at probe site $E=i$ and agrees well with the quadratic gap (Figure 3) near the exceptional point well. This can be seen by looking at the difference of the two gaps in Figure 6. On the other hand, while the Clifford linear gap agrees with the quadratic gap near the exceptional point (Figure 5), by itself it does not detect well the exceptional point (Figure 1). This is the motivation for defining the so called Clifford radial gap.
Figure 1:
$\bar{\mu}_{(0,E)}^{\text{C}}\left(X,H\right)$
Figure 2:
$\dot{\mu}_{(0,E)}^{\text{C}}\left(X,H\right)$
Figure 3:
$\mu_{(0,E)}^{\text{Q}}\left(X,H\right)$
Figure 4:
$\left|\bar{\mu}_{(0,E)}^{\text{C}}\left(X,H\right)-\dot{\mu}_{(0,E)}^{\text{C}}\left(X,H\right)\right|$
Figure 5:
$\left|\bar{\mu}_{(0,E)}^{\text{C}}\left(X,H\right)-\mu_{(0,E)}^{\text{Q}}\left(X,H\right)\right|$
Figure 6:
$\left|\dot{\mu}_{(0,E)}^{\text{C}}\left(X,H\right)-\mu_{(0,E)}^{\text{Q}}\left(X,H\right)\right|$
Table 1: Shown are the Clifford linear and radial gaps (Figure 1 and 2), quadratic gap (Figure 3) as well as the difference between the gaps in Figure 4 5, 6 of a Two Level System with $c=1$,$\Delta\gamma=2$ and position probe site $x=0$.
8 Applications to a line gapped system
Now we will provide an example of a physical system where the non-Hermiticity comes from a non-Hermitian Hamiltonian which is line gapped. Consider the generalized Haldane model over a bi-partite honeycomb lattice given in [haldane] with the following tight binding model hamiltonian;
$H$ $=\sum_{n_{A},n_{B}}\left((M-i\mu)|n_{A}\rangle\langle n_{A}|-(M+i\mu)|n_{B}\rangle\langle n_{B}|\right)-t\sum_{\langle n_{A},m_{B}\rangle}\left(|n_{A}\rangle\langle m_{B}|-|m_{B}\rangle\langle n_{A}|\right)$
$\qquad-t_{c}\sum_{\alpha=A,B}\sum_{\langle\langle n_{\alpha},m_{\alpha}\rangle\rangle}\left(e^{i\phi(n_{\alpha},m_{\alpha})}|n_{\alpha}\rangle\langle m_{\alpha}|+e^{-i\phi(n_{\alpha},m_{\alpha})}|m_{\alpha}\rangle\langle n_{\alpha}|\right).$
The first sum runs over all the lattice sites with $A$ and $B$ lattices having on-site energies ±$\pm M$. The second sum is kinetic energy with nearest neighbor coupling coefficient $t$. Lastly we have that the third sum is over next-nearest-neighbor pairs and has a direction-dependent phase factor $e^{i\phi(n_{\alpha},m_{\alpha})}$ a geometrically determined sign. When $\mu=0$ in the Hamiltonian, we have no lossy sites and the Hamiltonian is Hermitian. By making the model lossy, the Hamiltonian becomes non-Hermitian.
Our example tight binding model will be comprised of an inner topological insulator surrounded by a ring of trivial insulator which is then surrounded by a ring of a lossy trivial insulator, we call it the Haldane Heterostructure as in [cerjan_koekebier_Schulz_Baldes]. Again the loss comes from having on site energies being ±$\pm M-i\mu$ on the lossy Trivial insulator sites. The diagram for the tight binding model and unit cell is shown in Table 2. Furthermore, we will focus our attention on fixing the parameters for the tight binding model to the ones listed in Table 3.
Figure 7: Haldane Heterostructure Overview [Uncaptioned image] Figure 8: Haldane Hetero Structure Cell [Uncaptioned image]
Table 2: Shown in Figure 7 is the Haldane Heterostructure overview. Figure 8 shows the general cell for the Haldane Heterostructure
Layer $M$ $\mu$ $t$ $t_{c}$ $\phi$
Trivial Insulator (Lossy) $0.5\sqrt{3}$ $0.2$ $1$ 00 N/A
Trivial Insulator (Non-Lossy) $0.5\sqrt{3}$ 00 $1$ 00 N/A
Topological Insulator 00 00 $1$ $0.5$ $\frac{\pi}{2}$
Table 3: Parameters for Heterostructure
In both the spectral localizer and quadratic composite operator we will use $\kappa=0.5$. For the two positions operators we will take unit length between ‘A’ sites. They will end up being diagonal matrices as shown in Equation (15).
$X=\text{Diag}(x_{1},...,x_{i},...,x_{n}),\quad Y=\text{Diag}(y_{1},...,y_{n})$ (15)
The Hamiltonian will be non-Hermitian, and sparse for large system size. In some basis the diagonal entries will consist of complex (±$\pm M-i\mu$), purely real ±$\pm M$ and zeros as seen in Equation 16. In Equation 16, $C$ is a matrix with zeros on the diagonal.
$H=\text{Diag}(M-i\mu,...,-M-i\mu,...,M,...,-M,...,0,...,0)+C$ (16)
The 2d Clifford composite operator and the quadratic composite operator are as follows
$L^{\textup{2d}}_{(\kappa x,\kappa y,E)}(\kappa X,\kappa Y,H)$ $=\kappa(X-x)\otimes\sigma_{x}+\kappa(Y-y)\otimes\sigma_{y}+(H-E)\otimes\begin{bmatrix}1&\\ &0\end{bmatrix}+(H-E)^{\dagger}\otimes\begin{bmatrix}0&\\ &-1\end{bmatrix}$
$Q^{\textup{2d}}_{(\kappa x,\kappa y,E)}(\kappa X,\kappa Y,H)$ $=\kappa\left(X-x\right)^{2}+\kappa\left(Y-y\right)^{2}+(H-E)^{\dagger}(H-E).$
The pseudospectra tell us where we should expect to find approximate joint states with the corresponding imaginary probe site $E$ telling us how lossy the state is. We demonstrate via computation the Clifford linear and quadratic gap functions with varying position probe sites and fixed energy probe site in Table 4. In the same table we also demonstrate the differences between the various gap functions.
Figure 9:
$\bar{\mu}_{(\kappa\bm{x},0)}^{\text{C}}\left(\kappa\bm{X},H\right)$
Figure 10:
$\mu_{(\kappa\bm{x},0)}^{\text{Q}}\left(\kappa\bm{X},H\right)$
Figure 11:
$\left|\bar{\mu}_{(\kappa\bm{x},0)}^{\text{C}}\left(\kappa\bm{X},H\right)-\mu_{(\kappa\bm{x},0)}^{\text{Q}}\left(\kappa\bm{X},H\right)\right|$
Figure 12:
$\bar{\mu}_{(\kappa\bm{x},1)}^{\text{C}}\left(\kappa\bm{X},H\right)$
Figure 13:
$\mu_{(\kappa\bm{x},1)}^{\text{Q}}\left(\kappa\bm{X},H\right)$
Figure 14:
$\left|\bar{\mu}_{(\kappa\bm{x},1)}^{\text{C}}\left(\kappa\bm{X},H\right)-\mu_{(\bm{x},1)}^{\text{Q}}\left(\kappa\bm{X},H\right)\right|$
Figure 15:
$\bar{\mu}_{(\kappa\bm{x},i)}^{\text{C}}\left(\kappa\bm{X},H\right)$
Figure 16:
$\mu_{(\kappa\bm{x},i)}^{\text{Q}}\left(\kappa\bm{X},H\right)$
Figure 17:
$\left|\bar{\mu}_{(\kappa\bm{x},i)}^{\text{C}}\left(\kappa\bm{X},H\right)-\mu_{(\bm{x},i)}^{\text{Q}}\left(\kappa\bm{X},H\right)\right|$
Table 4: Shown are the Clifford linear gap (Figure 9,12, and 15), quadratic gap (Figure 10,13, and 16) as well as the difference between the gaps in Figure 11,14,and 17 of the Haldane Heterostructure with $(\bm{x},E)=(x,y,E)$ being the probe site with varying energy prove site $E=0,1,i$ along the rows of the table. Note that $(\bm{X},H)=(X,Y,H)$ are the two Position operators and Hamiltonian.
9 Conclusion
We have extended the theory of $\epsilon$-pseudospectrum [cerjan_loring_vides_2023] to allow for the study of systems that have more than one non-Hermitian operator. Along the way we have also proven some propositions about a special case of the non-Hermitian spectral localizer in [cerjan_koekebier_Schulz_Baldes] which we call the Clifford composite operator. We have also shown that even with perturbations to non-Hermitian $B_{i}$ the quadratic gap is local. With that said the benefits of the quadratic composite operator is that minimizing the corresponding gap function allows us to minimize eigen errors for various observables/matrices whithout increasing the computational complexity of the problem.
10 Acknowledgements
J.J.G. acknowledges support in part by an appointment to the Minority Educational Institution Student Partnership Program (MEISPP), sponsored by the U.S. Department of Energy and administered by the Oak Ridge Institute for Science and Education, and the Center for Integrated Nanotechnologies Summer Research Initiative. T.L. acknowledges support from the National Science Foundation, grant DMS-2349959. A.C. acknowledges support from the Laboratory Directed Research and Development program at Sandia National Laboratories. This work was performed, in part, at the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy (DOE) Office of Science. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for the U.S. DOE’s National Nuclear Security Administration under contract DE-NA-0003525. The views expressed in the article do not necessarily represent the views of the U.S. DOE or the United States Government.
[…]